OpenStudy (tparker22):

How many grams are in 1.68 x 1026 molecules of CO2? (molar mass=44.01 g/mol)

OpenStudy (sunnnystrong):

Okay: so Recall that 6.022 x 10^23 Molecules in one mole  

OpenStudy (geobit):

E=n*L E=1.68*10^26 L=6.02*10^23 By making "n" the subject, we get n=E/L When substituting the figures in thier corresponding variables and simplifying, we get n=279.06 mol Then from mole eqn. ( n= m/M) Where m is mass and M is molar mass, we can make "m" the subject to be "m=nM" By putting in the values and simplifying ,m (mass) =12281g.

OpenStudy (tparker22):

Thank you!!